### v1

# WildFireCan-MMD: A Multimodal dataset for Classification of User-generated Content During Wildfires in Canada

Abstract
Rapid information access is vital during wildfires, yet traditional data sources are slow and costly. Social media offers real-time updates, but extracting relevant insights remains a challenge. We present WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires, annotated across 13 key themes. Evaluating both Vision Language Models and custom-trained classifiers, we show that while zero-shot prompting offers quick deployment, even simple trained models outperform them when labelled data is available, by up to 23%. Our findings highlight the enduring importance of tailored datasets and task-specific training. Importantly, such datasets should be localized, as disaster response requirements vary across regions and contexts.
WildFireCan-MMD: A Multimodal dataset for Classification of User-generated Content During Wildfires in Canada
Braeden Sherritt Carleton University                        Isar Nejadgholi National Research Council Canada                        Marzieh Amini Carleton University
1 Introduction
Large-scale analysis of online interactions, using machine learning algorithms has enabled critical applications such as disaster response Fauzi (2023); Kumar et al. (2024b), mental health monitoringGhanadian et al. (2024); Dalal et al. (2024), detecting and combating hate speech in digital environments Kiritchenko et al. (2021); Kumar et al. (2024a), and sentiment and behavioral trend analysis Ma et al. (2024); Curto et al. (2024). Among these applications, analysis of social media data during natural disasters has gained increasing importance in recent years. Such data contains crucial details such as reports of casualties, infrastructure damage, urgent needs of affected individuals, and information about missing persons, among other critical information Houston et al. (2015); Alam et al. (2018). Using the shared information effectively can facilitate humanitarian aid efforts for more efficient management and faster recovery. This valuable data, however, is mixed in with noise, such as humorous content, advertisements, and other irrelevant information. Therefore, there is a critical need to develop machine learning models trained for sifting through this information.
To foster research in this area, Alam et al. (2018) have developed the CrisisMMD dataset to harness multimodal social media data for disaster response and humanitarian aid. This dataset was collected from Twitter and included earthquakes, hurricanes, wildfires, and floods occurring in 2017. However, the classes of CrisisMMD are generalized across all types of disasters, meaning they are not tailored to specific disaster scenarios. As a result, the classification scheme is broad, and certain nuances or unique information that may be relevant to a particular type of disaster could be overlooked. Various information, such as the nature of the damages and the type of assistance required, often differs depending on the specific disaster and its geographic context.
To address this limitation, we focus our work on gathering social media data specifically related to wildfires in Canada. Wildfires are becoming increasingly frequent and severe in this country due to the adverse effects of climate change. Although the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising about %12 of the total dataset and specific to the US context. Therefore, wildfires in Canada remain underexplored in social media analysis research. Upon inspecting our wildfire-specific social media dataset, we observed that existing categorization schemes do not capture the specificity of wildfire-related social media content. To address this, we devised thirteen tailored categories reflecting the unique information needs and communication patterns of Canadian wildfire contexts.
Then, we explored the feasibility of training multimodal classification models for triaging tweets during wildfires. In recent years, the research community has extensively worked on developing multimodal machine learning models for processing and categorizing social media data during natural disasters such as wildfires, hurricanes, and floods Ofli et al. (2020); Dwarakanath et al. (2021); Algiriyage et al. (2022); Basit et al. (2023); Koshy and Elango (2023); Hossain et al. (2022); Rezk et al. (2023); Khattar and Quadri (2022). Almost all of these works use the CrisisMMD dataset for their research, highlighting the scarcity of datasets on this topic.
To Benchmark the available models on WildFireCan-MMD, we evaluate the zero-shot classification capabilities of VLMs. Specifically, we prompt LLaVA (open-source) and gpt-4o-mini (proprietary) for label generation, and benchmark their performance against custom-trained multimodal classifiers using concatenative early and late fusion of RoBERTa and ViT modelsOfli et al. (2020); Basit et al. (2023). While VLMs offer a fast, label-free alternative, our results show that fine-tuned fusion models significantly outperform them. The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code is available at WildFireCanMMD-Github.  [^1] Please contact the authors for more information about data sharability.. The contributions of this study are summarized as follows:
  • We developed a 13-category taxonomy through multimodal topic modleing of wildfire-related user-generated content.
  • We introduced WildFireCan-MMD, the first Canada-specific multimodal wildfire dataset, comprising 4,688 posts from X in 2022–2024, labeled based on our taxonomy.
  • Uisng WildFireCan-MMD as benchamrk, we evaluate zero-shot VLMS on this task, finding that while closed-source models outperform open ones, both fall short without fine-tuning, which highlights the need for curated training data. Fine-tuned classifiers outperform zero-shot VLMs significantly; our best model achieves an F-score of 0.83, exceeding gpt-4o-mini by 23%.
2 Background and Literature Review
In this section, we review the literature on social media analysis for disaster response beyond the machine learning field, as well as the literature on the technical aspects of this project.
2.1 Social media analysis for disaster response
The growing interest among researchers in the role of social media during disasters has led to significant contributions in this field. Reference Fauzi (2023) conducted a bibliometric analysis to explore the use of social media in disaster management. This work provides a detailed review of the field, identifying key publications and future research directions, such as the implications of climate emergencies, integrating big data analytics, and managing knowledge in disaster contexts. Similarly, Muniz-Rodriguez et al. (2020) focused on the public health aspects of social media use during emergencies, noting its effectiveness in spreading emergency warnings and assessing needs after a disaster. Furthermore, Algiriyage et al. (2022) offered a systematic review of the application of deep learning in disaster response. They discuss the successes and challenges faced in this area, propose guidelines for future research, and consider the importance of using multimodal data in disaster scenarios. These contributions highlight the significant and varied ways social media can support disaster response efforts. Also, Acikara et al. (2023) explored how social media analytics can improve disaster response by enabling the rapid collection of time-sensitive information from affected areas. They noted the value of social media users as "citizen sensors" for quickly gathering essential data. Their review suggests that using social media analytics in disaster management, particularly for natural and human-made hazards, is a developing area that emphasizes the importance of human-centric insights, such as collective intelligence and location awareness. They also identify a gap in research on managing transport networks in real-time after disasters, which could enhance community resilience and sustainability. In other works, (Dwarakanath et al. (2021), Bukar et al. (2020), Mundottukandi et al. (2024)) discussed the crucial role of social media in managing crises, pointing out how platforms like Facebook and Twitter support communication and coordination among affected individuals and communities. Their reviews offer guidance for selecting research topics and aiding emergency teams in making informed decisions.
Our work is motivated by this literature and focuses on the use of machine learning models to triage social media data to potentially be used in all phases of disaster management.
2.2 Multimodal classification for disaster response
Expanding beyond text-based analysis of social media content, recent studies have emphasized multimodal analysis. We review this literature based on the available datasets and previously explored algorithms.
Datasets: Reference Dong et al. (2021) aimed to improve disaster relief by analyzing social media data to understand public attitudes and demand for relief supplies across different disaster types, using a textual Twitter dataset. They apply both qualitative and quantitative methods, including eight machine learning models, to assess public perception, offering insights for more effective disaster management. Also, Ochoa and Comes (2021) introduced a machine learning model that uses multimodal data, including satellite imagery and social media, to predict housing and shelter needs following disasters, demonstrating the model’s accuracy in forecasting such requirements. The lack of labeled social media imagery data for disaster response has been addressed by Alam et al. (2018), who created CrisisMMD, a comprehensive multimodal dataset from Twitter covering various natural disasters. This dataset, with thousands of manually annotated tweets and images from events like earthquakes, hurricanes, wildfires, and floods, provides valuable annotations for crisis management tasks. The dataset has been extensively utilized in research, shedding light on the complex communication patterns during disasters and the importance of integrating textual and visual data to better understand social media’s role in disaster contexts. The work by Ofli et al. (2020) split the problem into two distinct tasks. ’Task 1’ being informative or non-informative, to determine at a high level whether a given tweet is useful for humanitarian aid purposes. ’Task 2’ being the humanitarian categories or classes which the tweets are sorted into, for which they had a total of eight.
While previous works used a generalized scheme of annotations across various disasters, our work develops a wildfire-specific taxonomy of thirteen classes and introduces, WildFireCan-MMD, a Canadian dataset annotated with these labels.
Algorithms: Previous work by Ofli et al. (2020) introduced a joint representation learning from two parallel deep learning architectures where one architecture represents the text modality and the other represents the image modality. Image and text modalities are vectorized through VGG16 and Word2Vec+CNN architectures, respectively, followed by a dense layer for prediction. The work by Basit et al. (2023), expands on Ofli et al. (2020) by experimenting with multiple recent transformer structures, for both the text and image modalities, and compared them against the baseline reported in the work by Ofli et al. (2020). They found that using a combination of AlBERTLan et al. (2020) as a text embedder, and DeITTouvron et al. (2021) as an image embedder, provided the best results for the eight-class classification task. Reference Koshy and Elango (2023) implemented further experiments with text and image transformers and showed improvements by using bi-LSTMs instead of the traditional LSTM. Furthermore, they have noted that multiplicative fusion strategies produce better results on a classification task when combined with the previously described methods. Specifically, they demonstrated that using the fine-tuned RoBERTaLiu (2019) model for text, Vision Transformer model for image, biLSTM+attention mechanism with multiplicative fusion, for prediction, achieves the best results for the eight-class classification task described by Alam et al. (2018).
We build on the previous work and explore the existing structures on the newly annotated data, WildFireCan-MMDintroduced in this work.
2.3 Large Vision-Language Models
Recent advances in multimodal LLMs have enabled models to process image and text jointly, expanding their utility across a range of complex tasks. A comprehensive survey by Zhang et al. (2024) outlines the design, training strategies, and benchmark performance of 26 MM-LLMs, while Li et al. (2022) categorizes the evolution of VLMs into task-specific models, vision-language pretraining approaches, and those leveraging large-scale, weakly labeled image-text data. These VLMs now support zero- and few-shot generalization across diverse domains.
In this work, we evaluate gpt-4o-mini and LLaVA as zero-shot classifiers on the WildFireCan-MMD dataset. gpt-4o-mini, OpenAI’s most recent transformer-based VLM at the time Achiam et al. (2023), is closed-source and accessible only via API, limiting transparency around its architecture and training. In contrast, LLaVA Liu et al. (2023) is fully open-source and trained using GPT-4-generated responses paired with user-generated prompts and CLIP-based visual features Radford et al. (2021). It employs visual instruction tuning to align vision and language modalities and has shown strong performance on multimodal instruction-following benchmarks. The release of models like LLaVA reflects broader efforts to build high-performing, publicly accessible VLMs in response to the closed nature of commercial systems like GPT-4.
3 The WildFireCan-MMD Dataset
3.1 Data Collection
We used X’s Pro-Tier paid API to gather a diverse dataset of posts containing information, updates, and discussions related to the British Columbia/Alberta wildfires of 2022 and 2023, and the Jasper wildfires of 2024. This dataset contains tweets between the months of April and June for each year and consists of 4,688 unique image-text pairs. We collected this data using the Tweepy library and conducted a hashtag-based search, using hashtags we found to yield good results through experimentation. This was a crucial part of our data collection as we wanted to ensure we gathered tweets originating from Canada, that were also on topic. Due to not many users in Canada having their specific geo-location turned on for Twitter, we couldn’t simply scrape by region, or very little data would be returned. As a result, we had to adopt a hashtag-based search. Due to the uniqueness of the hashtags, which are only really used in Canada during wildfires, this yielded a dataset of relevant tweets with mostly users from the specified region. The dataset includes the timestamp of when the tweet was posted, the user-provided location, and, in some cases, the specific region or county from which the tweet originated. Table A.1 presents the full queries used for data collection.
3.2 Developing a Taxonomy
To examine the content of the tweets and discover the hidden, latent semantics of our data, we used the BERTopic to perform multimodal unsupervised topic modeling Grootendorst (2022); Falkenberg et al. (2022); Egger and Yu (2022); Medvecki et al. (2023). We then manually inspected the topics to identify the categories. In doing so, we considered the stakeholders who will be using this data. Implementation details of the BERTopic clustering are provided in Appendix B. In the end, we selected 13 classes which were important and recurrent enough to be considered categories for this dataset. The full list of categories and their descriptions is given in Table 1.
The selection of classes was guided by two primary considerations: the frequency of occurrence observed through manual annotation and the perceived relevance of classes to stakeholders in disaster scenarios. For example, ‘Warnings & Status Updates’, ‘Reports of Actions of Responders’, and ‘Infrastructure’ could be useful to first responders and emergency personnel. The ‘Evacuees,’General Information,’ and ‘Preparedness’ categories could be useful to the general public to be informed about the events happening around their locality. The ‘Political’ category could be useful to authorities to gauge public opinion on events surrounding the disaster.
Table 1: Categories of WildFireCan-MMD, their descriptions and the size of each class.
Category Description Class Size
Evacuees Information relating to evacuees, their movements, needs, location. 242
Smoke and Air Quality Tweets related to or showing signs of smoke or the current air quality. 1070
General Information General info about the wildfire situation, such as total hectares burned. 186
Preparedness Information for the general public to prepare themselves and their property. 265
Weather Reports Information relating to the weather with a specific location mentioned. 316
Warnings & Status Updates Fire bans in certain areas, new information about a specific area, updates from officials. 620
Reports of Actions of Responders Actions of responders within specific areas or times, including prescribed burns. 378
Infrastructure Detours, road closures, damage to infrastructure (e.g., utility poles, highways), repairs by crews. 293
Political Posts directed towards political figures or parties (excluding situation updates). 322
Support Mental health and financial support, temporary housing for livestock. 164
Insurance Information relating to insurance, employment insurance, and EI benefits. 142
Advertisement Posts about food, restaurants, off-topic ads for services (e.g., apps, air purifiers, not insurance-related). 116
Other No ’useful’ information, focusing on images (e.g., scenery), general complaining, or irrelevant content. 569
3.3 Data Labeling
We labeled the collected dataset according to the taxonomy of categories presented in Table 1. Labelling was done in-house by one of the authors, using Label Studio [^2] https://labelstud.io/. Since this task is not subjective, we did not consider multi-annotator labeling. Label Studio was chosen for its ability to search through the text fields of the dataset for specific keywords, rather than running through the dataset in sequence, which made labeling easier. For example, this allowed us to search for ’air quality’ and see all posts containing these keywords. This allowed us to focus on one category at a time and get the majority of a category labeled before moving on to another. When searching by keywords was no longer possible, labeling in sequence resumed.
Label: Reports of Actions of Responders
(a) Text: PAFD is on the scene of a fire in Roger Creek Park. There is no threat to structures at this time. #portalberni #BCWildfire
Label: Infrastructure
(b) Text: The wildfires in #AB have burned down this schoolhouse. #ABWildfires
Label: Warnings & Status Updates
(c) Text: Wildfire updates, July 24: Fire reaches Jasper townsite | ’Significant loss has occured’: Parks Canada | Firefighters working to protect critical infrastructure #abfire
Figure 1: Examples of labeled posts from WildFireCan-MMD.
Examples of labeled tweets are presented in Figure 1. Figure 2 displays the class distribution in WildFireCan-MMD. Also, for training and testing the custom classifiers, given in Section 4, we split the data into an 80/20 train/test split using stratified sampling, to ensure similar distributions of classes in train and test subsets. The size of labeled classes are presented in Table 1.
Figure 2: Class distribution of WildFireCan-MMD.
4 Classification Methodology
This section describes a range of algorithms evaluated for the 13-class classification of WildFireCan-MMD, utilizing text and image modalities. We also experimented with using Optical Character Recognition (OCR) to further extract more information from the images and create another channel of text to aid in classification. We experiment with using VLMs as zero-shot classifiers, and compare these results with fine-tuned transformer-based classifiers as well as to baseline models such as Random Forest, and Gaussian Naive Bayes. We present the implementation details of classification algorithms in the following.
4.1 VLM as Zero-Shot Classifiers
We crafted classification prompts for both gpt-4o-mini and LLaVa in an ’all-at-once’ method, meaning that with each query to the model, a list of all categories and their descriptions are given as a prompt, along with a sample (text + image) from the data to classify. We specify that the model’s task is to choose one of the provided categories and respond only with its corresponding letter. The system and user prompts can be seen in Table C.1.
4.1.1 gpt-4o-mini:
We used the available API for querying gpt-4o-mini 333version:gpt-4o-mini-2024-07-18. A specified content field in each query contains our system prompt, text from the Tweet, and user prompt. The image for the tweet is also sent with this payload.
The ’temperature’ is specified as being $0.1$ for reproducibility. ’max_tokens’ is set to be $1024$, as the text plus the image plus the model’s response will never be greater than this. The responses are saved as a JSONL file where each line contains the question ID from the questions file, as well as the model’s answer supplied as a single letter from the options shown in Table C.1. This then allowed us to calculate metrics such as balanced accuracy score and weighted precision, recall, and F1-score.
4.1.2 LLaVA:
As an open source model, querying LLaVA required a different approach than gpt-4o-mini. For our work, we opted for the 13B 4-bit version of LLaVa. The temperature was again set to $0.1$, and max_tokens set to $1024$. A custom conversation template was created using the original LLaVA codebase to construct our system prompt using the roles USER and ASSISTANT. The question file was created using the steps discussed above for gpt-4o-mini, the only difference being the addition of ’<>’ characters around each option as recommended by the source [^4] https://github.com/haotian-liu/LLaVA/blob/main/docs/Evaluation.md for querying LLaVA with a multiple-choice prompt. The text from the tweet is passed similarly to the method described above for gpt-4o-mini, and the image is converted to a tensor and passed with this to the model. The answers are saved similarly as described for gpt-4o-mini and saved, allowing us to calculate metrics such as balanced accuracy score and weighted precision, recall, and F1-score.
4.2 Custom Classifiers
To create a custom multimodal classifier for our dataset, the following steps were taken to implement concatenative early and late fusion.
Early and late fusion 2-head structure::This model consists of two architectures working in parallel, an image head and a text head. The image head uses the ViT transformer, and the text head uses the RoBERTa transformer. Another linear layer is added for prediction. Figure D.1 shows the structure of our text+image ’2-head’ multimodal classifier using concatenative early fusion. We created another model with the same strategy but utilizing late fusion instead, where each head has its own final linear layer and functions as an independent classifier. The two heads output probabilities, which are then averaged for a final result.
Training and Evaluation: A random seed was used for training reproducibility. Some notable hyperparameters used include a batch size of 16, an image size of ($384\times{384}\times{3}$) (as required by google/vit-base-patch16-384), and a max length for the text of 200. Twitter posts can have a maximum length of 200 characters, but after viewing the average and max length of all posts within the dataset, we learned that most users don’t go anywhere near this limit, and so this value could be decreased. We set the maximum number of epochs to be 20, and used Adam optimizer with a learning rate of $5e^{-5}$, and cross-entropy loss. Our experimentation found that the model converged at around 10-15 epochs. The model was then tested on the test set (9̃00 samples), and metrics were calculated, including balanced accuracy score and weighted precision, recall, and F1-score.
Additional Information with OCR: Trying to improve our classification metrics without gathering and labeling more data, we considered other ways in which we could extract more information from the data we had. Something we noticed in much of the data, specifically the Evacuees, Smoke & Air Quality, General Information, Preparedness, Warnings and Status Updates, and Weather Reports classes, was that many of the images contained valuable text which we believed could aid in classification. Although ViT and RoBERTA both use the transformer architecture, the way text is processed differs from images and doesn’t allow for ViT to differentiate text within images from the surrounding pixels in this current implementation.
This led us to consider ways in which we could extract text from the images in our dataset using the EasyOCR tool [^5] https://github.com/JaidedAI/EasyOCR. EasyOCR works by taking an image, preprocessing it, then applying the CRAFT algorithm to find bounding-box regions of text in an image. The pixels in these regions are sent to a text recognition model composed of a CRNN with three components: feature extraction, sequence labeling, and decoding to extract the text from these regions and turn them into strings as output. Deploying this model on our dataset, we extracted all of the text from every image and saved it to a new column as a pre-processing step. As a result, our dataset contains four columns: text, extracted_text, images, and labels. If an image was found to not have any text, a filler sentence was used: ’Image does not contain text.’.
Utilizing this extra column of text data, we added another RoBERTa head to our classifier. This new classifier, now with three separate heads concatenated and fed into a single linear layer, as appeared in Figure D.2. Note that this figure represents the early fusion version. A late fusion version was also created similarly as described for the 2-head classifier. The model was then trained with the same settings as described previously, for a similar number of epochs.
5 Experimental Results
Table 2 shows the calculated F1-scores for the VLMs, custom classifiers, and two baseline models per class and as an average over all classes. Also, figure E.1 shows the confusion matrices of the VLMs and the custom classifiers. Note that all metrics reported in these tables use ’weighted’ F1-scoring to better represent the class imbalance observed in our dataset.
Table 2: F1-scores per-class for VLMs, custom classifiers, and baseline models.
VLMs Custom Classifiers Baseline Classifiers
Class LLaVa gpt-4o-mini 2-head-early 3-head-early 2-head-late 3-head-late R. Forest Gaussian NB
Evacuees 0.17 0.37 0.77 0.73 0.79 0.74 0.29 0.50
General Information 0.14 0.16 0.59 0.46 0.53 0.59 0.35 0.67
Preparedness 0.61 0.65 0.82 0.81 0.86 0.84 0.52 0.80
Weather Reports 0.59 0.64 0.88 0.91 0.91 0.91 0.45 0.66
Warnings & Status Updates 0.36 0.45 0.80 0.76 0.77 0.74 0.50 0.57
Reports of Actions of Responders 0.29 0.63 0.89 0.86 0.85 0.83 0.48 0.74
Infrastructure 0.38 0.40 0.92 0.95 0.90 0.93 0.62 0.64
Political 0.70 0.77 0.84 0.84 0.84 0.85 0.28 0.78
Insurance 0.93 0.90 0.93 0.89 0.93 0.91 0.92 0.91
Advertisement 0.08 0.68 0.76 0.76 0.74 0.75 0.23 0.81
Smoke and Air Quality 0.76 0.79 0.86 0.86 0.86 0.85 0.58 0.70
Support 0.25 0.55 0.73 0.63 0.75 0.78 0.45 0.63
Other 0.05 0.54 0.78 0.78 0.76 0.79 0.36 0.60
Average 0.41 0.58 0.81 0.79 0.81 0.81 0.46 0.69
Weighted Average 0.45 0.60 0.83 0.81 0.82 0.82 0.48 0.67
LLaVa has limited capabilities.Testing on WildFireCan-MMD, We observed that LLaVa is not able to reach an acceptable result, performing close to the Random Forest baseline and underperforming gpt-4o-mini by a large margin. Specifically, we noted a large difference in F1-score between gpt-4o-mini and LLaVA for the Advertisement and Other classes (gpt4o-mini achieved between 0.5 and 0.6, whereas LLaVA reaches only less than 0.1). This can be explained by the fact that these two classes have the highest variance by far and are hard to handle for LLaVa. Specifically, the Other category is very diverse since tweets are labeled as so if they don’t fit into any of the other categories. Referring to confusion matrices (Figure E.1(a)), we can see that LLaVA struggles with labeling the Warnings & Status Updates class, where it often classified tweets from other classes as such, especially the Other class.
gpt-4o-mini shows strengths and limitations across various categories. On the other hand, gpt-4o-mini, being much larger and more sophisticated than LLaVA, outperforms it in almost every class except for the Insurance category, where it has been beaten by LLaVA by a couple of percent. Interestingly, gpt-4o-mini reaches a similar F1-score as custom classifiers on the Smoke & Air Quality, Advertisement, Insurance and Political classes, which combined encompass over a third of the dataset. This is probably due to the lower variance of these classes. On the contrary, the classes in which it suffers the most are Evacuees, General Information, Warnings & Status Updates, Infrastructure which contain some of the highest variance. It is difficult to construct a prompt, which encompasses the variance of these classes, while staying within the token limit. So, this may be one reason why gpt4-o performed worse than custom models in these categories. As seen in Figure E.1(b) we can see that gpt-4o-mini had the most trouble with the Warnings & Status Updates class, where it often confused it with the General Information and Evacuees classes. This effect caused poor performance in the Infrastructure class as well.
Custom classifiers are more robust and reliable. Referring to Table 2, we can see that the two and three-head classifiers provide similar results. Our 2-head classifiers have an average weighted F1-score of 0.83 for early fusion, and 0.82 for late fusion. The 3-head classifiers have an average weighted F1-score of 0.81 for early fusion, and 0.82 for late fusion. Surprisingly, the three-head classifier performs slightly worse in most classes than the two-head classifier, even though it receives extra contextual information from the images. One important consideration, however, when thinking about the utility and purpose of this work, is that the 3-head classifiers perform slightly better on key classes such as Infrastructure, Weather Reports, Smoke & Air Quality, General Information, and Support. Infrastructure is an important class because it can provide information to first responders and administrative personnel about key infrastructure that has been damaged or is currently in need of focus. Weather Reports is also important because it can provide insights and predictions about weather patterns and likely areas of new fires. General Information is useful to administrators and the general public as this class contains broad information about the situation as a whole.
6 Discussion
This work benchmarks VLMs on a new dataset. A critical consideration in evaluation of the VLMs is that WildFireCan-MMD establishes a novel benchmark for evaluating these models. Most evaluation research works are performed with pre-existing datasets, and these datasets are also highly likely to be used for training VLMs and do not represent a genuine test set. Our newly created dataset evaluates the real generalization capabilities of such models since it was not available to the creators of these models during training. This is important for information management during crises because new disasters have unique requirements.
The use of VLMs involves tradeoffs. Practical implications are involved in benchmarking with VLMs in a zero-shot setting. We discuss these implications from three perspectives: accuracy, cost and compliance with the responsible AI principles. The closed-source model, gpt-4o-mini, while being accurate to some extent, is expensive for large-scale deployment. Also, closed-source models often fail to comply with responsible AI principles due to a lack of transparency, privacy, and limited scrutiny and ethical oversight Hua et al. (2024). On the other hand, despite the free and open-source nature of LLaVA, its performance falls short in critical tasks such as disaster management. If performance is improved, the open-source nature of this model might make it suitable for disaster response, allowing compliance with responsible AI guidelines through offering access to the code base and ensuring ethical scrutiny. Particularly, the transparent nature of open source is crucial when handling sensitive data that may include personal information. Despite gpt-4o-mini performing better on our test set, it is important to keep these above considerations in mind. When labeled data is unavailable, VLMs can function as a zero-shot classifier, but in practice, neither model is particularly well suited to this specific type of task, emphasizing the necessity for a custom fine-tuned model.
Fine-tuning on task-specific data is essential. Results presented in Table 2 demonstrate that even simple approaches such as Random Forest and Gaussian NB, when trained with task-specific labelled data, provide similar or better results than what is achieved from zero-shot VLMs. Even though our dataset is modest in size, it clearly showcases the benefits of developing a labeled dataset and fine-tuning a custom classifier. Further, our results show that high degrees of accuracy can be achieved for this task by fine-tuning advanced classification structures, such as pre-trained transformer-based text and image encoders.
The custom classifiers we have developed serve as a foundational step for triaging tweets during wildfires, providing an initial performance benchmark and demonstrating our approach’s feasibility. However, to enhance the accuracy and robustness of our model, future work will need to adopt a more comprehensive approach to experimentation. One area for improvement is the application of different fusion strategies, such as multiplicative fusion, which has been shown to better capture the relationship between different modalities Koshy and Elango (2023).
7 Conclusion
General-purpose VLMs, while powerful, struggle to perform reliably in high-stakes, context-rich scenarios like disaster response. To address this gap, we present WildFireCan-MMD, a wildfire-specific, multimodal dataset of social media posts (text + image) curated from X, accompanied by a domain-focused taxonomy derived through multimodal topic modeling and manual annotation. Unlike prior disaster datasets, ours is tailored exclusively to wildfire contexts, making it a valuable benchmark for evaluating true generalization in VLMs. As climate-induced wildfires grow in frequency and complexity, this dataset supports the development of robust, transparent, and adaptive AI tools for detection, coordination, and recovery.
Limitations
While our study introduces a novel wildfire-specific multimodal dataset and demonstrates the effectiveness of fine-tuned task-specific models, several limitations should be acknowledged.
The fisrt limitation is the scope and size of the dataset. Although WildFireCan-MMD provides a targeted and realistic benchmark for wildfire-related content, its size remains modest compared to large-scale datasets used to train or evaluate general-purpose models. This limits the diversity of linguistic and visual expressions captured and may affect the generalizability of our results to other wildfire events or regions.
Platform and geographic bias might be present in our data. The dataset is sourced exclusively from the X platform and focuses on Canadian wildfires. This introduces both platform-specific and regional biases in language use, imagery, and public discourse. As a result, models trained on this dataset may not perform as well when applied to other social platforms or international wildfire contexts.
Moreover, zero-shot evaluation of VLMs introduces limitations. Prompt engineering for zero-shot setups is a heuristic process, and better prompts might yield different results.
Appendix A Data collection Queries
Table A.1 shows the full queries used to collect data from X.
Year Query
2022/2023 (#BCwildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire) -has:videos has:images lang:en -is:retweet -is:quote -is:reply
2024 (#BCwildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire OR #JasperStrong OR #JasperWildfire OR #JasperAB) -has:videos has:images lang:en -is:retweet -is:quote -is:reply
Table A.1: The full queries used for scraping data from X.
Appendix B Topic modeling parameters
In order to optimize the performance of BERTopic on our dataset, careful consideration of hyperparameters is key McInnes et al. (2018). First, text and image inputs are vectorized using the pre-trained “Clip-ViT-B-32” [^6] https://huggingface.co/sentence-transformers/clip-ViT-B-32 model. Then, BERTopic uses the UMAP algorithm for dimension reduction, as it keeps a significant portion of the high-dimensional local structure in lower dimensionality. We use the default parameter settings of BERTopic, which reduces the dimensionality to 5, while keeping the size of the local neighborhood at 15. After dimension reduction, BERTopic uses HDBSCAN to cluster the vectors. HDBSCAN is a density-based algorithm that does not force data points to clusters, as it can consider them outliers. HDBSCAN has been shown to be effective when paired with UMAP, since UMAP maintains the local structure in the lower-dimensional space.
After forming the clusters, BERTopic extracts keywords from each topic to uncover their central theme. For that, we set the parameters so that stop words are removed and both unigrams and bigrams are considered as candidates of topic keywords.
Employing BERTopic on our dataset, we first let the algorithm determine the number of latent topics instead of setting a set number of topics to sort the data into, as at this stage, we were unsure of what sort of topics were contained in our data. BERTopic had sorted our data into over 100 fine-grained topics. This was still relatively large to manage, and it was difficult to determine a humanitarian category for each fine-grained topic. After some experimentation, we systematically decreased the total number of topics to determine which topics remained over time. Once we had reduced our topics to around 15, it became clearer which topics were the strongest in our dataset, and we were able to begin to see the humanitarian utility of some of these topics. Some examples of useful topics were posts about smoke and air quality, weather reports, and reports of actions of first responders. We also noticed some less useful posts, such as advertisements and political debates/complaints. Some examples of these topics, as well as their corresponding keywords, can be seen in Figure B.1.
Using BERTopic in the way described proved to be useful, as it gave us a good understanding from the beginning of what our future categories were to look like. Relying on this method alone was not enough, however, and having to look through the data manually at some point was inevitable. Many rounds of manual annotation ensued on a small portion of the data, after having it randomly mixed, in order to avoid constructing our categories on themes that may only exist at a certain time during the wildfire event. Through performing these steps, keeping in mind our discovered topics through BERTopic and humanitarian goals, new categories were added or dropped based on what was seen in the data. A final round of labelling then commenced, and the entire dataset was manually labeled.
Keywords: [crews, safe, poles, damaged, continue, lines, working restore, power, conditions, restore]
(a) Useful
Keywords: [highway, closed, closures, abroads, detour, road, route, bchwy4, cameron, cameron lake]
(b) Useful
Keywords: [’yycnow’, ’yycbusiness’, ’yycbusiness yycnow’, ’yyc yycbusiness’, ’abfires abwildfire’, ’calgary’, ’startups’, ’wordify’, ’smallbusiness’, ’smallbusiness startups’]
(c) Non-useful
Figure B.1: Examples of useful/non-useful topics from topic modeling on our dataset.
Appendix C Prompts for zero-shot classification
We show the full prompt used to evaluate VLMs in Table C.1.
Table C.1: System and User prompt used for gpt-4o-mini and LLaVa.
System Prompt:
"You are an assistant who is being given an image and text pair as a Twitter post which was created during a natural disaster event. Your task is to use information from both the text and the image to decide which option the post should be labeled as. You must pay close attention to each option when deciding which label to use."
User Prompt:
"Which option should this post be labeled as?
A. Evacuees (information relating to evacuees, their movements, needs, location, etc)
B. General Information (GENERAL facts about the wildfire situation, hectares burned)
C. Preparedness (information for the general public to prepare themselves and property for wildfires)
D. Weather Reports (information relating to the weather, satellite imagery, radar imagery)
E. Warnings & Status Updates (warnings/updates to the public from authoritative bodies, fire bans, specific information relating to a certain time or area)
F. Reports of Actions of Responders (prescribed burns, responders responding to a specific location)
G. Infrastructure (road closures, damaged buildings or property, traffic)
H. Political (mentions of political or public figures or parties)
I. Insurance (mentions of insurance)
J. Advertisement (information about restaurants, food, apps or services)
K. Smoke & Air Quality (information about smoke or air quality, masks, breathing)
L. Support (information about financial, mental health, or other types of support for people)
M. Other (the post does not fit well in one of the previous categories)
You may only answer with the chosen option’s letter."
Appendix D Structures of custom classifiers
Figures D.1 and D.2 show the structures of the finetuned models. We described these models in detail in Section 4.2.
Figure D.1: Custom multimodal 2-head early fusion classifier structure.
Figure D.2: Custom multimodal 3-head early fusion classifier structure.
Appendix E Confusion matrix for all classifiers
For deeper insights into classifiers’ performances, we show the confusion matrices of all classifiers in Figure E.1.
a) LLaVA, b) gpt-4o-mini, c) 2head-early, d) 3head-early, e) 2head-late, f) 3head-late
Figure E.1: Confusion matrices of classifiers and VLMs.