### v2

# WildFireCan-MMD: A Multimodal Dataset for Classification of User-Generated Content During Wildfires in Canada

Abstract
Rapid information access is vital during wildfires, yet traditional data sources are slow and costly. Social media offers real-time updates, but extracting relevant insights remains a challenge. We present WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires, annotated across twelve key themes. Evaluating both vision-language models and custom-trained classifiers, we show that while zero-shot prompting offers quick deployment, even simple trained models outperform them when labelled data is available. Our best-performing transformer-based fine-tuned model reaches 83% f-score, outperforming gpt4 by 23%. Also, we demonstrate how this model can be used to uncover trends during wildfires. Our findings highlight the enduring importance of tailored datasets and task-specific training. Importantly, such datasets should be localized, as disaster response requirements vary across regions and contexts.
WildFireCan-MMD: A Multimodal Dataset for Classification of User-Generated Content During Wildfires in Canada
Braeden Sherritt Carleton University                        Isar Nejadgholi National Research Council Canada                        Marzieh Amini Carleton University
1 Introduction
Large-scale analysis of online interactions, using machine learning algorithms has enabled critical applications such as disaster response Fauzi (2023); Kumar et al. (2024b), mental health monitoring Ghanadian et al. (2024); Dalal et al. (2024), detecting and combating hate speech in digital environments Kiritchenko et al. (2021); Kumar et al. (2024a), pandemic discourse analysis Imran et al. (2022) and sentiment and behavioral trend analysis Ma et al. (2024); Curto et al. (2024). Among these applications, analysis of social media data during natural disasters has gained increasing importance in recent years Alam et al. (2018); Ofli et al. (2020); Alam et al. (2021); Dwarakanath et al. (2021); Algiriyage et al. (2022); Hossain et al. (2022); Khattar and Quadri (2022); Basit et al. (2023); Koshy and Elango (2023); Rezk et al. (2023). Such data contains crucial details such as reports of casualties, infrastructure damage, and urgent needs of affected individuals, among other critical information. The effective use of the shared information can facilitate humanitarian aid efforts for more efficient management and faster recovery Houston et al. (2015). This valuable data, however, is mixed in with noise, such as humorous content, advertisements, and other irrelevant information. Therefore, there is a critical need to develop an automated method for sifting through this information.
To foster research in this area, Alam et al. (2018, 2021) developed two multimodal datasets, CrisisMMD and HumAID, both collected from Twitter, including several disasters such as earthquakes and hurricanes. However, the classes of these datasets are generalized across all types of disasters, meaning they are not tailored to specific disaster scenarios. As a result, the classification scheme is broad, and certain nuances or unique information that may be relevant to a particular type of disaster could be overlooked. Also, while the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising about %12 of the total dataset and specific to the US context. The HumAID dataset by Alam et al. (2021) also contains wildfire data from Canada, but only a couple of thousand samples from 2016.
Therefore, wildfires in Canada remain underexplored in social media analysis research. To address this limitation, we focus our work on gathering social media data specifically related to Canadian wildfires, as they are becoming increasingly frequent and severe due to the adverse effects of climate change.
Upon initial inspection of our Canadian wildfire-specific social media dataset, we observed that existing categorization schemes do not capture the specificity of this data. To address this, we devised a novel taxonomy of twelve categories reflecting the unique information needs and communication patterns of Canadian wildfire contexts. Using our developed taxonomy, we manually labeled 4,688 wildfire-related tweets from Canada and created WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires.
To benchmark the available classification models on WildFireCan-MMD, we evaluated the zero-shot classification capabilities of several VLMs, as well as the feasibility of training custom classifiers. Our best-performing classifier is a fine-tuned model, which uses RoBERTa Liu (2019) and Vision Transformer (ViT) models Ofli et al. (2020); Basit et al. (2023) to jointly process text and image parts of tweets. While WildFireCan-MMD is sourced from data posted in 2022-2024, we collected an additional set of unlabeled tweets across seven years (2018-2025). We then used our best-performing classifier to analyze this data and uncover trends and patterns of social media use in Canada. The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code, is available at WildFireCanMMD-Github [^1] Please contact the authors for more information about data sharability., under CC-BY-NC license. The contributions of this study are summarized as follows:
  • We devised a twelve-category taxonomy through multimodal topic modeling of Canadian wildfire-related user-generated content, and introduced WildFireCan-MMD, the first Canada-specific multimodal wildfire dataset, comprising 4,688 posts from X in 2022–2024, manually labeled based on our taxonomy.
  • We compared custom multimodal transformer models trained on WildFireCan-MMD with several VLMs and classical methods. Fine-tuned classifiers outperform zero-shot VLMs significantly; our best model achieves an F-score of 0.83, exceeding gpt-4o-mini by 23%.
  • We collected an additional 46,279 posts from X, spanning across seven years, used our trained classifier to label them, and reported the uncovered trends and patterns of social media use during wildfires in Canada.
2 Background and Literature Review
2.1 Social media analysis for disaster response
Significant research has been conducted on the role of social media during disasters. Fauzi (2023) conducted a bibliometric analysis of social media use in disaster response and offered a comprehensive overview of key studies and emerging research directions. Acikara et al. (2023) highlighted the potential of social media analytics in disaster response by leveraging users as “citizen sensors” to rapidly gather critical, time-sensitive information. Muniz-Rodriguez et al. (2020) focused on the public health aspects of social media use during emergencies, noting its effectiveness in spreading emergency warnings and assessing needs after a disaster. In other works, Dwarakanath et al. (2021), Bukar et al. (2020), and Mundottukandi et al. (2024) discussed the crucial role of social media in managing crises, pointing out how platforms like Facebook and Twitter support communication and coordination among affected individuals and communities. Furthermore, in a systematic review of the application of deep learning in disaster response, Algiriyage et al. (2022) discussed the successes and challenges faced in this area, proposed guidelines for future research, and highlighted the importance of using multimodal data in disaster scenarios.
Our work is motivated by this literature and focuses on using machine learning models to triage Canada-specific multimodal social media data to inform disaster response and management.
2.2 Multimodal classification for disaster response
We review this literature based on the available datasets and previously explored algorithms.
Datasets: While many works utilize textual datasets sourced from social media to classify posts during disasters Olteanu et al. (2014); Dong et al. (2021); Ochoa and Comes (2021), it is now generally considered that multimodal approaches offer stronger results Ofli et al. (2020); Abavisani et al. (2020); Ramirez et al. (2025).
Alam et al. (2018) addressed the shortage of multimodal labeled social media data for disaster response by introducing CrisisMMD, a widely used dataset containing annotated tweets and images from various natural disasters. Later extended by datasets like HumAID Alam et al. (2021), CrisisMMD has become a foundational resource for studying disaster-related communication patterns and the combined use of text and images. It remains the most prominent dataset in this domain, used in many recent studies Basit et al. (2023); Koshy and Elango (2023); Shetty et al. (2024).
While previous works such as CrisisMMD used a generalized scheme of annotations across various types of different disasters, our work develops a wildfire-specific taxonomy of twelve categories and introduces WildFireCan-MMD, a new Canadian wildfire dataset.
Algorithms: Previous work by Ofli et al. (2020) introduced a joint representation learning from two parallel deep learning architectures where one architecture represents the text modality and the other represents the image modality. Image and text modalities are vectorized through VGG16 and Word2Vec+CNN architectures, respectively, followed by a dense layer for prediction. The work by Basit et al. (2023), expands on Ofli et al. (2020) by experimenting with multiple recent transformer structures, for both the text and image modalities, and compared them against the baseline reported in the previous work. They found that using a combination of AlBERT Lan et al. (2020) as a text embedder, and DeIT Touvron et al. (2021) as an image embedder, provided the best results for the eight-class classification task. Koshy and Elango (2023) implemented further experiments with text and image transformers and showed improvements by using bi-LSTMs instead of LSTMs. Furthermore, they have noted that multiplicative fusion strategies produce better results on a classification task when combined with the previously described methods. Specifically, they demonstrated that using the fine-tuned RoBERTa Liu (2019) model for text, Vision Transformer model for image, biLSTM+attention mechanism with multiplicative fusion, for prediction, achieves the best results for the eight-class classification task described by Alam et al. (2018). We build on the previous work and explore the existing structures on the newly annotated data, WildFireCan-MMD.
In addition to fine-tuned structures, we explore using VLMs in a zero-shot setting Zhang et al. (2024); Li et al. (2022). Specifically, we evaluate three VLMs, GPT-4 Achiam et al. (2023), LLaVA Liu et al. (2023), and SmolVLM Marafioti et al. (2025), to assess performance across a range of capabilities and computational requirements. GPT-4 serves as a state-of-the-art, closed-source benchmark, while LLaVA and SmolVLM are open-source models that use visual instruction tuning and are optimized for varying levels of hardware, reflecting ongoing efforts to build accessible, high-performing multimodal systems.
3 The WildFireCan-MMD Dataset
3.1 Data Collection
To construct our dataset, we utilized X’s Pro-Tier paid API to collect a diverse set of posts surrounding the British Columbia and Alberta wildfires of 2022 and 2023, as well as the Jasper wildfires of 2024. The dataset comprises 4,688 unique image-text pairs gathered from tweets posted between April and June of each respective year. Data collection was conducted using a hashtag-based search strategy. Table A.1 presents the full queries used for data collection. To curate the hashtags, we started with seed hashtags and used an iterative process of adding new hashtags by collecting sample tweets. This approach was necessary due to the limited availability of geotagged content, as most users on the platform do not enable precise geo-location. Instead, we identified and utilized hashtags that, through iterative experimentation, were found to be effective indicators of relevant content originating from Canada during wildfire events. The specificity of these hashtags ensured that the collected tweets were not only on-topic but also predominantly from users in the targeted geographic regions. The dataset includes the timestamp of when the tweet was posted, the user-provided location, and, in some cases, the specific region or county from which the tweet originated.
3.2 Taxonomy Development
To examine the content of the tweets and discover the hidden latent semantics of our data, we used BERTopic to perform multimodal unsupervised topic modeling Grootendorst (2022); Falkenberg et al. (2022); Egger and Yu (2022); Medvecki et al. (2023). Further explanation of the BERTopic clustering method are provided in Appendix B.
Employing BERTopic on our dataset, we first let the algorithm determine the number of latent topics instead of setting number of topics to sort the data into. When the number of clusters were unspecified, BERTopic had created over 100 fine-grained topics. This was relatively large to manage, and difficult to determine a humanitarian category for each fine-grained topic. We gradually decreased the number of topics, and for each number, we manually inspected the representative visual samples of topics and keywords generated by BERTopic. Once we had reduced our topics to around 15, it became clearer which topics were the strongest in our dataset, and we were able to begin to see the humanitarian utility of some of these topics. Some examples of useful topics were posts about infrastructure, weather reports, and reports of actions of first responders. We also noticed some less useful posts, such as advertisements and political debates/complaints. Some examples of these topics, including representative images and corresponding keywords, generated by BERTopic, can be seen in Figure B.1. We then manually inspected the topics and selected 12 classes that were informative and consistently appeared across a varying maximum number of clusters to be considered designated categories for this dataset.
The full list of categories and their descriptions is given in Table 1. We also include an “Other” to accommodate the examples that do not fit into any of these 12 categories. The selection of classes was guided by two primary considerations: the frequency of occurrence observed through manual annotation and the perceived relevance of classes to stakeholders in disaster scenarios. For example, ‘Warnings & Status Updates’, ‘Reports of Actions of Responders’, and ‘Infrastructure’ could be useful to first responders and emergency personnel. The ‘Evacuees,’General Information,’ and ‘Preparedness’ categories could be useful to the general public to be informed about the events happening around their locality. The ‘Political’ category could be useful to the media to gauge public opinion and policies on events surrounding the disaster.
Table 1: Categories of WildFireCan-MMD, their descriptions and total count of each class.
Category Description Count
Evacuees Information relating to evacuees, their movements, needs, location. 242
Smoke and Air Quality Tweets related to or showing signs of smoke or the current air quality. 1070
General Information General info about the wildfire situation, such as total hectares burned. 186
Preparedness Information for the general public to prepare themselves and their property. 265
Weather Reports Information relating to the weather with a specific location mentioned. 316
Warnings & Status Updates Fire bans in certain areas, new information about a specific area, updates from officials. 620
Reports of Actions of Responders Actions of responders within specific areas or times, including prescribed burns. 378
Infrastructure Detours, road closures, damage to infrastructure (e.g., utility poles, highways), repairs by crews. 293
Political Posts directed towards political figures or parties (excluding situation updates). 322
Support Mental health and financial support, temporary housing for livestock. 164
Insurance Information relating to insurance, employment insurance, and EI benefits. 142
Advertisement Posts about food, restaurants, off-topic ads for services (e.g., apps, air purifiers, not insurance-related). 116
Other No ’useful’ information, focusing on images (e.g., scenery), general complaining, or irrelevant content. 569
3.3 Data Labeling
We labeled the collected dataset according to the taxonomy of categories presented in Table 1. Over the course of labeling, we ensured that the data did not contain personal information. Labeling was carried out in-house by one of the authors using Label Studio [^2] https://labelstud.io/. To ensure the reliability of the labeling process, we first discussed the definitions and boundaries of each category among the research team and reached a shared understanding. Through this discussion, we determined that the task was not largely subjective, as the categories are well-defined and grounded in observable cues (e.g., specific topics, hazards, or references in the text or image). Therefore, we did not consider multi-annotator labeling necessary. Examples of labeled tweets are presented in Figure 1. The size of labeled classes are presented in Table 1.
Label: Reports of Actions of Responders
(a) Text: PAFD is on the scene of a fire in Roger Creek Park. There is no threat to structures at this time. #portalberni #BCWildfire
Label: Infrastructure
(b) Text: The wildfires in #AB have burned down this schoolhouse. #ABWildfires
Label: Warnings & Status Updates
(c) Text: Wildfire updates, July 24: Fire reaches Jasper townsite | ’Significant loss has occured’: Parks Canada | Firefighters working to protect critical infrastructure #abfire
Figure 1: Examples of labeled posts from WildFireCan-MMD.
4 Classification Methodology
This section describes a range of algorithms evaluated for the 13-class (12 categories of the taxonomy +“Other”) classification of WildFireCan-MMD. In general, we employ VLMs as zero-shot classifiers, and compare the results with fine-tuned transformer-based classifiers as well as to classical methods such as decision tree, and Gaussian naive bayes, as minimum baselines.
Since each post on X can have up to 4 associated images, we duplicated the text and paired with each image as an individual sample. For training and testing the custom classifiers, we split the data into an 80/20 train/test split using stratified sampling, to ensure similar distributions of classes in train and test subsets. We use the same test set for all of the classifiers, including zero-shot VLMs.
4.1 VLMs as Zero-Shot Classifiers:
We considered three VLMs, GPT-4 Achiam et al. (2023), LLaVA Liu et al. (2023), and SmolVLM Marafioti et al. (2025) and tested them in a zero-shot setting. We created classification prompts for all three VLMs where, with each query to the model, a list of all categories and their descriptions is given as a prompt along with a data instance (text + image) to classify. We specify that the model’s task is to choose one of the provided categories and respond only with its corresponding letter. The system and user prompts can be seen in Table C.1.
gpt-4o-mini: We used the available API for querying gpt-4o-mini 333version:gpt-4o-mini-2024-07-18. A specified content field in each query contains our system prompt, text and image from the tweet, and user prompt. The responses are saved as a file where each line contains the question ID from the questions file, as well as the model’s answer supplied as a single letter from the options shown in Table C.1. This then allows for the calculation of metrics.
LLaVA: We opted for the 13B 4-bit version of LLaVA due to hardware constraints. Using a combination of a publicly available Colab notebook  [^4] https://github.com/camenduru/LLaVA-colab and the original code from the open-source LLaVA repository  [^5] https://github.com/haotian-liu/LLaVA, the model was loaded onto the GPU within Colab. The answers are saved similarly as described for gpt-4o-mini.
smolVLM: Being the smallest of the VLMs experimented with, only requiring <10GB of GPU ram, experiments were able to be performed locally without quantization. Due to its inclusion in the Transformers library, loading and querying the model with the previously created prompts was simpler and cheaper than the other methods, while taking a similar amount of time to process.
4.2 Custom Classifiers
Prior work by Koshy and Elango (2023) demonstrated that a multimodal architecture combining RoBERTa with a BiLSTM and ViT, using multiplicative fusion and an additional attention mechanism, achieved the best performance on the CrisisMMD dataset. We evaluated this architecture, among several others, on our WildFireCA-MMD dataset. Our findings indicate that the BiLSTM and additional attention components did not improve performance in our context. Instead, a simpler architecture employing early fusion via feature concatenation of RoBERTa and ViT embeddings yielded the best results. We experimented with several model structures described in the following.
2-head structure:This model consists of two parallel architectures, an image head and a text head. The image head uses the ViT666’google/vit-base-patch16-384’ transformer, and the text head uses the RoBERTa777’roberta-base’ transformer. A linear layer is added for prediction. Figure D.1 shows the structure of our text+image ’2-head’ multimodal classifier using concatenative early fusion.
3-head structure: To enhance classification without more manual labeling, we extracted text from images using EasyOCR [^8] https://github.com/JaidedAI/EasyOCR. The extracted text was saved in a new column (extracted_text), defaulting to a filler if none was found. Our classifier was updated to include a third RoBERTa head for this column, as shown in Figure D.2.
Training and Evaluation: A random seed was used for reproducibility. Some notable hyperparameters used include a batch size of 16, an image size of ($384\times{384}\times{3}$) (as required by our ViT model). We set the maximum number of epochs to be 20, and used Adam optimizer with a learning rate of $5e^{-5}$, and cross-entropy loss. Our experimentation found that the model converged at around 10-15 epochs. The model was then tested on the test set (~1000 samples), and metrics were calculated, including balanced accuracy score and weighted precision, recall, and F1-score.
4.3 Baseline Classifiers
To establish baseline, we evaluated several classic classifiers on our dataset, namely Decision Tree, Gaussian Naive Bayes, and K-Nearest Neighbors (KNN). For textual data, we extracted TF-IDF feature matrices by applying the TfidfVectorizer with max_features=5000. The image data, was resized to 64$\times$64 pixels. The resulting TF-IDF and image matrices were concatenated and used as input to the classifiers. Hyperparameters were set to default values in Scikit-learn  [^9] https://scikit-learn.org/stable/.
5 Experimental Results
Table 2 shows the calculated F1-scores for the VLMs, custom classifiers, and baseline models per class and as an average over all classes. Also, Figures E.1 and E.2 show the confusion matrices of the custom classifiers and VLMs, respectively. Note that all metrics reported in these tables use ’weighted’ F1-scoring to better represent the class imbalance observed in our dataset.
Some classes are harder to distinguish than others. In general, we observe a large variance of classification F-scores across classes. For example, all classifiers struggle with classifying examples of Evacuees and General Information. We explain this observation by the great variability of information in these classes. On the other hand, classes with less information variety, such as Insurance and Smoke and Air Quality, are more manageable for most classifiers to detect.
Table 2: F1-scores per-class for VLMs, custom and baseline classifiers. Maximum results are in bold.
VLMs Custom Classifiers Baseline Classifiers
Class LLaVA GPT-4o-mini smolVLM 2-head 3-head Dec. Tree Gauss. NB KNN
Evacuees 0.17 0.37 0.25 0.77 0.73 0.47 0.53 0.29
General Information 0.14 0.16 0.00 0.59 0.46 0.23 0.46 0.21
Preparedness 0.61 0.65 0.23 0.82 0.81 0.57 0.72 0.49
Weather Reports 0.59 0.64 0.00 0.88 0.91 0.46 0.63 0.52
Warnings & Status Updates 0.36 0.45 0.00 0.80 0.76 0.45 0.59 0.43
Reports of Actions of Responders 0.29 0.63 0.00 0.89 0.86 0.45 0.70 0.40
Infrastructure 0.38 0.40 0.00 0.92 0.95 0.47 0.75 0.52
Political 0.70 0.77 0.00 0.84 0.84 0.52 0.75 0.31
Insurance 0.93 0.90 0.92 0.93 0.89 0.93 0.93 0.89
Advertisement 0.08 0.68 0.00 0.76 0.76 0.22 0.65 0.40
Smoke and Air Quality 0.76 0.79 0.50 0.86 0.86 0.60 0.69 0.51
Support 0.25 0.55 0.06 0.73 0.63 0.30 0.60 0.33
Other 0.05 0.54 0.27 0.78 0.78 0.35 0.67 0.33
Weighted Average 0.45 0.60 0.20 0.83 0.81 0.48 0.67 0.43
VLMs struggle with nuanced disaster communication, underperforming compared to classic classifiers. Table 2 shows that VLMs, including LLaVA, GPT-4o-mini, and smolVLM, consistently underperform across most disaster-related classes. Their weighted F1-scores range from 0.20 (smolVLM) to 0.60 (GPT-4o-mini), all falling short of the best-performing baseline (Gaussian Naive Bayes, 0.67). Similarly, the Decision Tree classifier achieved a weighted score of 0.48, comparable to LLaVA’s 0.45. The performance of smolVLM in particular was notably poor, with a weighted average of only 0.20 and zero scores in multiple categories. These results suggest that, despite their general-purpose capabilities, current VLMs struggle to perform reliably in domain-specific, fine-grained classification tasks, such as the case for WildFireCan-MMD.
Custom classifiers are more robust and reliable. In contrast to VLMs, our custom classifiers (2-head and 3-head architectures) demonstrate superior and more stable performance. With weighted F1-scores of 0.83 and 0.81, they outperform all baselines and VLMs across nearly all categories. The 2-head classifier achieves top scores in Insurance (0.93), Reports of Actions of Responders (0.89), Smoke and Air Quality (0.86). The 3-head model benefits from incorporating OCR-extracted text, performing particularly well in Infrastructure (0.95), and Weather reports (0.91). While incorporating OCR-extracted textual information doesn’t improve the classification results overall, it might be a favorable approach when certain classes are deemed to be more critical for certain response strategies.
6 Analysis of Trends in Data
After labeling our dataset and determining our best-performing classifier (the 2-head classifier described in section 4.3 and Figure D.1), we proceeded to collect another 46,279 posts from X. The collection methodology was the same as explained in section 3.1, where we gathered tweets using a curated set of wildfire-related hashtags specific to each year, selected based on known wildfire events and hotspot locations across affected regions. We combined this with a keyword-based search, then combined the results and removed duplicates. The full set of queries used for collecting this new data can be seen in Appendix F.
First, we extracted the author locations from the samples. Our goal was to obtain the province-level region where each sample is coming from. After extracting the author locations, which could be either automatically generated or user-provided, we ran the queries through OSMNX [^10] https://osmnx.readthedocs.io/en/stable/. This allowed us to automatically search for the locations of cities or towns that users provided and extract the province level information. If a users location was not found, or not in Canada, it was labeled as such. The distribution of posts by province can be seen in Figure 2, showing that the majority of data is posted from British Columbia and Alberta. Additional information about this data, including a word cloud of the text and the frequency distribution of the posts by word count can be found in Figures 1(a) and 1(b).
Figure 2: Distribution of collected tweets based on author location.
Using our best-performing classifier with an F1-score of 83%, we obtained predictions for the new data. This automated classification approach allowed for efficiency in processing the large volume of newly collected data without having to manually label each individual sample. This of course, will not uphold the same gold standard of labeling that a human annotator could achieve, and manual labeling will have to be done in the future. Individual classifications should be approached with skepticism, but this method should still allow for general trends in the data to be analyzed on a broader scale.
Figure 3 (left) shows the total number of posts per week sorted by year. This demonstrates a significant event occurred the week of June 3, 2023, and July 22, 2024, due to several major fires in the country Reuters Staff (2023); CBC News (2024). Figure 3 (right) illustrates weekly trends in predicted labels during 2023. Sharp spikes in Smoke and Air Quality and Warnings & Status Updates likely correspond to major wildfire events occurring in Alberta and BC. Warnings & Status Updates and Evacuees follow similar patterns around the week of Aug 12, which corresponds to the state of emergency declared as the city of West Kelowna was evacuated. Overall, at least 35,000 people were under evacuation order and another 30,000 under evacuation alert, as of August 19 Williams (2023). Reports of Actions of Responders and Support also rise during these peaks, reflecting coordinated emergency efforts.
Figure 3: Weekly post distributions: (left) overall yearly trend, (right) class-wise distribution in 2023.
7 Discussion
This work benchmarks VLMs on a new dataset. A critical consideration in evaluation of the VLMs is that WildFireCan-MMD establishes a novel benchmark for evaluating these models. Most evaluation research works are performed with pre-existing datasets, and these datasets are also highly likely to be used for training VLMs and do not represent a genuine test set. Our newly created dataset evaluates the real generalization capabilities of such models since it was not available to the creators of these models during training. This is important for information management during crises because new disasters have unique requirements.
The use of VLMs involves tradeoffs. Practical implications are involved in benchmarking with VLMs in a zero-shot setting. We discuss these implications from three perspectives: accuracy, cost and compliance with the responsible AI principles. The closed-source model, gpt-4o-mini, while being accurate to some extent, is expensive for large-scale deployment. Also, closed-source models often fail to comply with responsible AI principles due to a lack of transparency, privacy, and limited scrutiny and ethical oversight Hua et al. (2024). On the other hand, despite the free and open-source nature of LLaVA, its performance falls short in critical tasks such as disaster management. Also, smolVLM, while being both open-source and computationally affordable, has a performance far below acceptable rates. Therefore, when labeled data is unavailable, VLMs might be considered as a zero-shot classifier, but in practice, these models are not particularly well suited to this specific type of task, emphasizing the necessity for a custom fine-tuned model.
Fine-tuning on task-specific data is essential. Results presented in Table 2 demonstrate that even simple approaches such as decision tree and Gaussian NB, when trained with task-specific labeled data, provide similar or better results than what is achieved from zero-shot VLMs. Even though our dataset is modest in size, it clearly showcases the benefits of developing a labeled dataset and fine-tuning a custom classifier. Further, our results show that high degrees of accuracy can be achieved for this task by fine-tuning common classification structures, such as pre-trained transformer-based text and image encoders. These findings align with other works such as Nadeem et al. (2025).
Custom classifiers can be used to discover social media use patterns. The primary function of the custom classifier is to perform real-time labeling of social media data during wildfire events, enabling the timely extraction of actionable information to support disaster response efforts. Beyond real-time deployment, our post-hoc analysis of a large unlabeled dataset illustrates the classifier’s capacity to identify latent patterns and temporal trends, providing a scalable tool for retrospective analysis and evidence-based policy development.
To enhance the accuracy and robustness of our model, future work can consider the application of different fusion strategies, such as multiplicative fusion, which has been shown to better capture the relationship between different modalities Koshy and Elango (2023). Fine-tuning of VLMs is also a promising avenue of improving the task which should be explored in future works Duan et al. (2024); Zhai et al. (2024).
8 Conclusion
Despite the advancements in general-purpose VLMs, they struggle to perform reliably in high-stakes, context-rich scenarios like disaster response. This low accuracy, alongside the resource-intensive nature of such models, necessitates developing smaller fine-tuned models, specialized for specific regions and contexts. To address this gap, we present WildFireCan-MMD, a wildfire-specific, multimodal dataset of social media posts curated from X, accompanied by a domain-focused taxonomy derived through multimodal topic modeling and manual annotation. Unlike prior disaster datasets, ours is tailored exclusively to wildfire contexts, making it a valuable benchmark for evaluating true generalization in VLMs, and suitable for training specialized models. As climate-induced wildfires grow in frequency and complexity, this dataset supports the development of robust, transparent, and adaptive AI tools for detection, coordination, and recovery.
9 Acknowledgment
This work was funded by the National Program Office of the National Research Council Canada, under the New Beginning Initiative Program.
Limitations
While our study introduces a novel wildfire-specific multimodal dataset and demonstrates the effectiveness of fine-tuned task-specific models, several limitations should be acknowledged.
The first limitation is the scope and size of the dataset. Although WildFireCan-MMD provides a targeted and realistic benchmark for wildfire-related content, its size remains modest compared to large-scale datasets used to train or evaluate deep learning models. This limits the diversity of linguistic and visual expressions captured and may affect the generalizability of our results to other wildfire events or regions.
Platform and geographic bias might be present in our data. The dataset is sourced exclusively from the X platform and focuses on Canadian wildfires. This introduces both platform-specific and regional biases in language use, imagery, and public discourse. As a result, models trained on this dataset may not perform as well when applied to other social platforms or international wildfire contexts.
Moreover, zero-shot evaluation of VLMs introduces limitations. Prompt engineering for zero-shot setups is a heuristic process, and better prompts might yield different results.
Finally, there is a broader risk of misapplication. Deploying these models in real-world disaster contexts without appropriate oversight, interpretability mechanisms, or human-in-the-loop verification may lead to misclassification, misinformation amplification, or missed signals, potentially affecting emergency coordination and public trust.
Appendix A Data collection Queries
Table A.1 shows the full queries used to collect the WildFireCan-MMD data from X.
Year Query
2022/2023 (#BCwildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire) -has:videos has:images lang:en -is:retweet -is:quote -is:reply
2024 (#BCwildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire OR #JasperStrong OR #JasperWildfire OR #JasperAB) -has:videos has:images lang:en -is:retweet -is:quote -is:reply
Table A.1: The queries used for collecting the data for WildFireCan-MMD.
Appendix B Topic modeling parameters
In order to optimize the performance of BERTopic, careful consideration of hyperparameters is key McInnes et al. (2018). First, text and image inputs are vectorized using the pre-trained “Clip-ViT-B-32” [^11] https://huggingface.co/sentence-transformers/clip-ViT-B-32 model. Then, BERTopic uses the UMAP algorithm for dimension reduction, as it keeps a significant portion of the high-dimensional local structure in lower dimensionality. We use the default parameter settings of BERTopic, which reduces the dimensionality to 5, while keeping the size of the local neighborhood at 15. After dimension reduction, BERTopic uses HDBSCAN to cluster the vectors. HDBSCAN is a density-based algorithm that does not force data points to clusters, as it can consider them outliers. HDBSCAN has been shown to be effective when paired with UMAP, since UMAP maintains the local structure in the lower-dimensional space.
After forming the clusters, BERTopic extracts keywords from each topic to uncover their central theme. For that, we set the parameters so that stop words are removed and both unigrams and bigrams are considered as candidates of topic keywords.
Once we had reduced our topics to around 15, it became clearer which topics were the strongest in our dataset, and we were able to begin to see the humanitarian utility of some of these topics. Some examples of useful topics were posts about smoke and air quality, weather reports, and reports of actions of first responders. We also noticed some less useful posts, such as advertisements and political debates/complaints. Some examples of these topics, as well as their corresponding keywords, can be seen in Figure B.1.
Keywords: [crews, safe, poles, damaged, continue, lines, working restore, power, conditions, restore]
(a) Useful
Keywords: [highway, closed, closures, abroads, detour, road, route, bchwy4, cameron, cameron lake]
(b) Useful
Keywords: [’yycnow’, ’yycbusiness’, ’yycbusiness yycnow’, ’yyc yycbusiness’, ’abfires abwildfire’, ’calgary’, ’startups’, ’wordify’, ’smallbusiness’, ’smallbusiness startups’]
(c) Non-useful
Figure B.1: Examples of useful/non-useful topics from topic modeling on our dataset.
Appendix C Prompts for zero-shot classification
We show the full prompt used to evaluate the aforementioned VLMs in Table C.1. The following hyperparameters are used for all VLMs: temperature=0.1, max_new_tokens=1024.
Table C.1: System and User prompt used for VLMs.
System Prompt:
"You are an assistant who is being given an image and text pair as a Twitter post, which was created during a natural disaster event. Your task is to use information from both the text and the image to decide which option the post should be labeled as. You must pay close attention to each option when deciding which label to use."
User Prompt:
"Which option should this post be labeled as?
A. Evacuees (information relating to evacuees, their movements, needs, location, etc)
B. General Information (GENERAL facts about the wildfire situation, hectares burned)
C. Preparedness (information for the general public to prepare themselves and property for wildfires)
D. Weather Reports (information relating to the weather, satellite imagery, radar imagery)
E. Warnings & Status Updates (warnings/updates to the public from authoritative bodies, fire bans, specific information relating to a certain time or area)
F. Reports of Actions of Responders (prescribed burns, responders responding to a specific location)
G. Infrastructure (road closures, damaged buildings or property, traffic)
H. Political (mentions of political or public figures or parties)
I. Insurance (mentions of insurance)
J. Advertisement (information about restaurants, food, apps or services)
K. Smoke & Air Quality (information about smoke or air quality, masks, breathing)
L. Support (information about financial, mental health, or other types of support for people)
M. Other (the post does not fit well in one of the previous categories)
You may only answer with the chosen option’s letter."
Appendix D Structures of custom classifiers
Figures D.1 and D.2 show the structures of the finetuned models. We described these models in detail in Section 4.3.
Figure D.1: Custom multimodal 2-head early fusion classifier structure.
Figure D.2: Custom multimodal 3-head early fusion classifier structure.
Appendix E Confusion matrices of classifiers
For deeper insights into classifiers’ performances, we show the confusion matrices of all classifiers in Figures E.1 and E.2.
(a) 2-head classifier.
(b) 3-head classifier.
Figure E.1: Confusion matrices of custom classifiers.
(a) LLaVA.
(b) GPT-4.
(c) SmolVLM.
Figure E.2: Confusion matrices of VLMs.
Appendix F Queries Used for Additional Data Collection :
Below is the list of queries and keywords used to collect additional unlabeled data discussed in Section 6.
2018
Hashtag Query: (#BCwildfire OR #britishcolumbiawildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire)
Keyword Query: (alberta OR british columbia OR Prince George OR Grande Praire OR Waterton OR Bulkley Nechako OR Nadina Lake OR Kootenay OR Crowsnest Pass OR Medicine Lake OR Comstock Lake OR Tugwell Creek OR Sooke OR Nanaimo Lakes OR Tweedsmuir OR Johnny Creek OR Alkali Lake OR Lutz Creek OR Shovel Lake OR Nadina Lake OR Verdun Mountain OR Silver Lake OR Tommy Lakes OR Island Lake OR Chutanli Lake) (wildfire OR forest fire)
2019
Hashtag Query: (#ABWildfire OR #albertawildfire OR #ABFire)
Keyword Query: (alberta OR calgary OR edson OR Fort McMurray OR Grande Prairie OR High Level OR Lac La Biche OR Whitecourt OR Steen River OR Chuckegg Creek OR Peace River OR Slave Lake OR Wood Buffalo National Park) (wildfire OR forest fire)
2020
Hashtag Query: (#BCwildfire OR #britishcolumbiawildfire OR #BCfire OR #ABwildfire OR #albertawildfire OR #ABFire OR #SKwildfire OR #sasksatchewanwildfire OR #SKfire OR #YTwildfire OR #yukonwildfire OR #YTfire OR #NTwildfire OR #northwestterritorieswildfire OR #NTfire OR #NWTwildfire OR #NWTfire OR #MBwildfire OR #manitobawildfire OR #MBfire OR #ONwildfire OR #ontariowildfire OR #QCwildfire OR #quebecwildfire OR #QCfire)
Keyword Query: (british columbia OR alberta OR sasksatchewan OR yukon OR northwest territories OR manitoba OR ontario OR quebec) (wildfire OR forest fire)
2021
Hashtag Query: (#MBwildfire OR #manitobawildfire OR #MBfire OR #ONwildfire OR #ontariowildfire OR #SKwildfire OR #sasksatchewanwildfire OR #SKfire OR #pafire OR #ontariofire OR #manitobafire OR #sasksatchewanfire)
Keyword Query: (manitoba OR ontario OR sasksatchewan OR british columbia) (wildfire OR forest fire)
2022
Hashtag Query: (#YTwildfire OR #yukonwildfire OR #YTfire OR #yukonforestfire OR #NTwildfire OR #northwestterritorieswildfire OR #NTfire OR #NWTwildfire OR #NWTfire OR #nwtforestfire)
Keyword Query: (#yzf OR #nwt OR #Yellowknife OR Yukon OR Northwest Territories OR Whitehorse OR Yellowknife OR Dawson City OR Great Slave Lake OR Norman Wells OR Inuvik OR Watson Lake OR Hay River OR Fort Smith OR Tuktoyaktuk OR Behchoko) (wildfire OR forest fire) (-alaska -Eielson -CityofNorthPole)
2023
Hashtag Query: (#BCwildfire OR #britishcolumbiawildfire OR #BCfire OR #ABwildfire OR #albertawildfire OR #ABFire OR #SKwildfire OR #sasksatchewanwildfire OR #SKfire OR #YTwildfire OR #yukonwildfire OR #YTfire OR #NTwildfire OR #northwestterritorieswildfire OR #NTfire OR #NWTwildfire OR #NWTfire OR #MBwildfire OR #manitobawildfire OR #MBfire OR #ONwildfire OR #ontariowildfire OR #QCwildfire OR #quebecwildfire OR #QCfire OR #CanadaOnFire OR #CanadaWildfire OR #CanadaFires OR #CanadaIsOnFire)
Keyword Query: (ontario OR quebec OR sasksatchewan OR british columbia OR manitoba OR northwest territories OR yukon OR alberta) (wildfire OR forest fire)
2024
Hashtag Query: (#JasperStrong OR #JasperWildfire OR #JasperAB OR #BCwildfire OR #britishcolumbiawildfire OR #BCfire OR #ABwildfire OR #albertawildfire OR #ABFire OR #CanadaOnFire OR #CanadaWildfire OR #CanadaFires OR #CanadaIsOnFire)
Keyword Query: (canada OR ontario OR quebec OR sasksatchewan OR manitoba OR northwest territories OR yukon OR british columbia OR alberta OR jasper) (wildfire OR forest fire)
Note: Every query above includes the suffix -has:videos has:images lang:en -is:retweet -is:quote -is:reply to filter for original English-language posts with images only.
Appendix G Unlabeled dataset
A word cloud of the text data of the unlabeled dataset collected in Section 6, and also the frequency distribution of the posts by word count, are shown in Figures 1(a) and 1(b).
(a) Word cloud of the posts.
(b) Frquency distribution of the posts w.r.t their lengths.
Figure G.1: Word cloud of the textual part of the unlabeled tweets and the frequency distribution of them regarding their word counts.